Not algebra but I am really stuck, please help
Q1. which of the points A(-1,2) B(2,13) C (0,3) lie on the line with the equation y=5x + 3? What is the gradient of the line?

Q2.
2a gh
__ x ___
fg 4

Answer by venugopalramana(3286)

Not algebra but I am really stuck, please help
Q1. which of the points A(-1,2) B(2,13) C (0,3) lie on the line with the equation y=5x + 3? What is the gradient of the line?
IT MEANS YOU HAVE TO PUT X AND Y VALUES IN THE GIVEN EQN. AND SEE WHETHER
LHS = RHS OR NOT
IF YES..THE POINT LIES ON THE LINE
IF NO..THE POINT DOES NOT LIE ON THE LINE
A...X=-1..Y=2....LHS = Y = 2....RHS = 5X+3 =5*-1+3=-5 +3 =-2...NO
B...X=2...Y = 13 ..LHS = Y =13...RHS..= 5X +3 = 5*2 +3 =13 ..OK..YES
C...X=0...Y=3........LHS = Y = 3....RHS = 5*0+3 = 3....OK...YES


Q2.
2a gh
__ x ___
fg 4